# Explicit estimates for the logarithmic derivative and the reciprocal of the Riemann zeta-function

### v3

@@ 6 Concluding Remarks
Here we state some possible ways to improve our estimates. The most impactful would be having a larger zero-free region or a higher Riemann height. Using an asymptotically stronger zero-free region is also helpful. See for instance [7], where a Littlewood-style zero-free region is used to obtain bounds of order $\log t/\log\log t$. Minor improvements can be made by a more careful treatment of the sum over zeros and the integral arising from the application of Lemma 1. Other possibilities are: an improvement on the Borel–Carathéodory by specialising it more to this problem, proving a uniform (in $\sigma$) bound for $\arg\zeta(s)$, or using stronger estimates for $\zeta(s)$ under $RH$ or $RH(T)$.
@@
It is evident from the results that the closer $s$ is to the boundary of the zero-free region, and hence a potential zero, the larger the constants. However, the large constants can also be attributed in part to the factor $(R+r)/(R-r)$ in Theorem 1, which increases rapidly as $r$ approaches $R$. This contributes significantly to the overall estimate, even if $|f(s_{0})|$ is small, and seems to suggest that Theorem 1 in its current form might not be optimal for this problem.
The main advantage of our method using Lemma 1 is that it eliminates the need for Theorem 1 in estimating $|1/\zeta(s)|$, and also reduces the number of applications of the theorem when estimating $|\zeta^{\prime}(s)/\zeta(s)|$. It might be possible to remove this entirely, by directly bounding the real and imaginary part of $|\zeta^{\prime}(s)/\zeta(s)|$, as one can easily prove an analogue of Lemma 1 which relates $\textup{Im}(f^{\prime}/f)$ to $\arg f$. To use such a result in the case of $f(s)=\zeta(s)$, besides dealing with the corresponding sum over zeros, a uniform bound (over a range of $\sigma$) on $S_{\sigma}(t):=(1/\pi)\arg\zeta(\sigma+it)$ is needed.
@@
To get unconditional results we have (due to knowledge of the Riemann height) assumed a partial Riemann hypothesis to prove some of our results. As an example, for $W=5.5587$, the condition (37) in Theorem 5 is only fulfilled for $t\geq 3.22\cdot 10^{9}$. The conditions on $d$ (which depends on $t$) are what prevents us from taking $W$ too small, and this is where the Riemann height and $RH(T)$ helps us. In fact we only need a ‘weak’ version of $RH$, for instance in (26), that up to height $T$, no zeros lie to the right of $1/2-R^{\prime}$. The downside to this is that we have not used better available bounds that require the full strength of $RH$. For example, under $RH$ we could have used that $\zeta(1/2+it)=O(\exp(C\log t/\log\log t))$ for some $C>0$ to prove something stronger (see [26, §14], [24], [23], and [6] for more details).
@@
Even in the unconditional results used, there is room for improvement. As an illustration, Corollary 5, though asymptotically superior, beats Corollary 4 only when $t$ is large. In the case of $\sigma=1$, this happens for $t\geq\exp(161975)$. This is in part due to the relatively large constants in Corollary 5, and a direct way of lowering these constants is an improvement in the bound $|\zeta(1+it)|$, which correspondingly lowers $C_{3}$ in Lemma 6.
@@
Finally, we conclude with a discussion on the trigonometric inequality (24) used in many of our above proofs, in particular when bounding $|1/\zeta(s)|$ for $\sigma>1$. Under certain conditions, this method is superior to using the trivial bound. In general, when studying zero-free regions for $\zeta(s)$, increasing the degree of the non-negative trigonometric polynomial improves the results one gets. Surprisingly, the same cannot be said in our setting (with the exception of results in Section 5.4), and the classical polynomial we used is actually optimal. Take for instance Lemma 12, which is used in Theorems 4 and 5. Here the overall order of the bound is $O(\log t)$, and in this scenario the article [19] has an in-depth discussion on why the classical polynomial is best.
@@
On the other hand, in Theorems 6 and 7, the question is whether higher degree polynomials lead to a reduction in the exponent $11/12$, even if overall constants get worse. This is an interesting question to consider, since in this case the answer is less obvious, although preliminary calculations suggest that the classical polynomial is still the optimal one.
@@ Acknowledgements
We thank Tim Trudgian for his guidance throughout the writing of this paper. We are also very grateful to Bryce Kerr, Olivier Ramaré, and Aleksander Simonič for the helpful discussions.

### v4

@@ 6 Concluding Remarks
Here we state some possible ways to improve our estimates. The most impactful would be having a larger zero-free region or a higher Riemann height. Using an asymptotically stronger zero-free region is also helpful. See for instance [7], where a Littlewood-style zero-free region is used to obtain bounds of order $\log t/\log\log t$. Minor improvements can be made by a more careful treatment of the sum over zeros and the integral arising from the application of Lemma 2. For Theorem 4, a better smoothing argument could be used in place of (45).
Other possibilities for improvement are: a refinement on the Borel–Carathéodory by specialising it more to this problem, using stronger estimates for $\zeta(s)$ under $RH$ or $RH(T)$, or obtaining a good bound on ${\textup{Re}}(\zeta^{\prime}(s)/\zeta(s))$ for use in (42).
@@
It is evident from the results that the closer $s$ is to the boundary of the zero-free region, and hence a potential zero, the larger the constants. However, the large constants can also be attributed in part to the factor (R+r)/(Rr)(R+r)/(R-r) in Theorem 1, which increases rapidly as $r$ approaches $R$. This contributes significantly to the overall estimate, even if $|f(s_{0})|$ is small, and seems to suggest that Theorem 1 in its current form might not be optimal for bounding $|\zeta^{\prime}(s)/\zeta(s)|$.
An advantage of obtaining a bound on ${\textup{Re}}(\zeta^{\prime}(s)/\zeta(s))$ is to reduce the usage of Theorem 1, when bounding $|1/\zeta(s)|$. This is because we would be able to apply such a bound directly to (42), whereas we currently use ${\textup{Re}}(\zeta^{\prime}(s)/\zeta(s))\leq|\zeta^{\prime}(s)/\zeta(s)|$. This alternative could be approached via Lemma 2, provided one were able to deal with the resulting sum over zeros well.
@@
To get unconditional results we have (due to knowledge of the Riemann height) assumed a partial Riemann hypothesis to prove some of our results. The downside to this is that we have not used better available bounds that require the full strength of $RH$. For example, under $RH$ we could have used that $\zeta(1/2+it)=O(\exp(C\log t/\log\log t))$ for some $C>0$ to prove something stronger (see [26, §14], [25], [24], and [6] for more details).
@@
Even in the unconditional results used, there is room for improvement. As an illustration, Corollary 5, though asymptotically superior, beats Corollary 4 only when $t$ is large. In the case of $\sigma=1$, this happens for $t\geq\exp(2.2\cdot 10^{5})$. This is in part due to the relatively large constants in Corollary 5, and a direct way of lowering these constants is an improvement in the bound $|\zeta(1+it)|$, which correspondingly lowers $C_{3}$ in Lemma 6.
@@
Finally, we conclude with a discussion on the trigonometric inequality (20) used in many of our above proofs, in particular when bounding $|1/\zeta(s)|$ for $\sigma>1$. Under certain conditions, this method is superior to using the trivial bound. In general, when studying zero-free regions for $\zeta(s)$, increasing the degree of the non-negative trigonometric polynomial improves the results one gets. Surprisingly, the same cannot be said in our setting, and the classical polynomial we used is actually optimal. Take for instance Lemma 12, which is used in Theorem 4. Here the overall order of the bound is $O(\log t)$, and in this scenario the article [20] has an in-depth discussion on why the classical polynomial is best.
@@
On the other hand, in Theorem 5, the question is whether higher degree polynomials lead to a reduction in the exponent $11/12$, even if overall constants get worse. This is an interesting question to consider, since in this case the answer is less obvious, although preliminary calculations suggest that the classical polynomial is still the optimal one.
@@ Acknowledgements
We thank Tim Trudgian for his guidance throughout the writing of this paper. We also thank the referee for the care taken in going through our article. Finally, we are very grateful to Andrew Fiori, Bryce Kerr, Olivier Ramaré, Aleksander Simonič, for the helpful discussions, and especially to Harald Helfgott for pointing out an error in an earlier version of this paper.